# Characterizing higher Auslander(-Gorenstein) Algebras

### v1

@@ 1.7. Cotorsion Pairs
Let now $\mathcal{X}$ and $\mathcal{Y}$ be full subcategories of the abelian category $\mathcal{A}$. Recall that a pair $(\mathcal{X},\mathcal{Y})$ is called a cotorsion pair (or cotorsion theory) if $\mathcal{X}={}^{\perp_{1}}\mathcal{Y}$ and $\mathcal{Y}=\mathcal{X}^{\perp_{1}}$, see for instance, [EJ11, Definition 7.1.2]. In this case, the class $\mathcal{X}$ is called a cotorsion class and the class $\mathcal{Y}$ is called a cotorsion-free class.
@@
Salce’s lemma (which goes back to [Sa79]) asserts that $(i)$ and $(ii)$ are equivalent in the case where $\mathcal{A}$ is a module category, see, e.g. [EJ11, Proposition 7.1.7]. The proof of this lemma shows that if the abelian category $\mathcal{A}$ has enough injectives, then $(i)\Rightarrow(ii)$ and if $\mathcal{A}$ has enough projectives, then $(ii)\Rightarrow(i)$.
In contrast to the case with torsion pairs, the constructions $x,\tilde{x}$ and $y,\tilde{y}$ are not functorial. But they define functors from $\mathcal{A}$ to $\frac{\mathcal{X}}{\mathcal{X}\cap\mathcal{Y}}$ and $\frac{\mathcal{Y}}{\mathcal{X}\cap\mathcal{Y}}$, respectively [BBOS, Lemma 2.11].
Note that there are slightly different definitions of cotorsion pairs which have appeared in the literature, see [BR07, KS03]. By [BR07, Remark V.3.2], there is no difference between the notion of a complete cotorsion pair that we recalled it in the above definition and the notion of a cotorsion pair in the sense of Beligiannis-Reiten [BR07, Definition V.3.1]. Also, cotorsion pairs in the sense of Krause-Solberg [KS03, Definition 2.1] are hereditary complete cotorsion pairs.
@@ 1.8. Projectivization and Injectivization
restricts to an equivalence $\mathrm{add}(M)\simeq\mathrm{proj}(\Gamma)$; see, for instance, [ARS, Proposition II.2.1 (c)].
@@
The first statement is well known; see, for instance, [ARS, Proposition II.2.5].
For the second statement, since $P$ is finitely generated, $\mathrm{Hom}_{\Lambda}(P,-)$ commutes with direct sums, thus it establishes an equivalence between $\mathrm{Add}(P)$ and ${\mathrm{Proj}}(\Gamma)$. Then the proof proceeds as that of [ARS, Proposition II.2.5]. ∎
@@
(a) This is the first lemma of [Au71, Section 4].
@@
The two functors $\mathbf{D}\mathrm{Hom}_{\Gamma}(-,\mathcal{Q})$ and $\mathrm{Hom}_{\Gamma}(I,-)$ are exact, in particular, right exact, so by Watts’ theorem (see, for example, [Ro09, Corollary 5.47]), they are isomorphic to a tensor functor. It suffices to check that $\mathbf{D}\mathrm{Hom}_{\Gamma}(\Gamma,\mathcal{Q})\cong\mathbf{D}\mathcal{Q}\cong\mathrm{Hom}_{\Gamma}(I,\Gamma)$. In fact, since $\mathbf{D}\mathcal{Q}$ is finitely generated projective, its double $\Gamma$-dual $\mathrm{Hom}_{\Gamma}(\mathrm{Hom}_{\Gamma^{{\rm{op}}}}(\mathbf{D}\mathcal{Q},\Gamma),\Gamma)=\mathrm{Hom}_{\Gamma}(I,\Gamma)$ is isomorphic to itself.
@@
commutes with direct sums, so the equivalence $\mathrm{add}(\mathcal{Q})\simeq\mathrm{inj}(\Lambda)$ extends to $\mathrm{Add}(\mathcal{Q})\simeq{\mathrm{Inj}}(\Lambda)$. As in the proof of the first lemma of [Au71, Section 4], the latter equivalence extends to another equivalence
@@
1.9. Injective Resolutions
For the convenience of the reader, this subsection is devoted to state some important results about injective resolutions which we’ll make use of them in several proofs.
Recall that an Iwanaga-Gorenstein algebra is a two-sided Noetherian ring such that its injective dimensions both as a left and as a right $\Gamma$-module are finite, which are equal by Zaks [Za69, Lemma A] ( see also [EJ11, Proposition 9.1.8]).
Theorem 1.7 (Iwanaga [Iw79, Theorem 2]).
Let $\Gamma$ be an Iwanaga-Gorenstein algebra with the minimal injective resolution
$0\longrightarrow{{}_{\Gamma}\Gamma}\longrightarrow I^{0}\longrightarrow I^{1}\longrightarrow\cdots.$
Then $\bigoplus_{i\geq 0}I^{i}$ is an injective cogenerator.
Theorem 1.8 (Iwanaga [Iw80, Theorem 3 (1)]).
Let $\Gamma$ be an Artinian Iwanaga-Gorenstein algebra with the minimal injective resolution
$0\longrightarrow{{}_{\Gamma}\Gamma}\longrightarrow I^{0}\longrightarrow\cdots\longrightarrow I^{n}\longrightarrow 0$
Then the projective dimension of every direct summand $I$ of $I^{n}$ is $n$, i.e. ${\rm{pd}}({}_{\Gamma}I)=n$.
Hence, for Gorenstein algebras of Gorenstein dimension $n+1$, we have the following proposition.
@@ 1.9. Injective Resolutions
If $M\in\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)$, then $M=\bigoplus_{l=1}^{r}M_{l}$, where $M_{l}$ is an injective indecomposable $\Gamma$-module and of projective dimension at most $n$. By Theorem 1.7, each $M_{l}$ is a direct summand of $I^{j}$ with $10\leq j\leq n+$ 1. But ${\rm{pd}}(M_{l})\leq n$, so by Theorem 1.8, $M_{l}$ is not a direct summand of $I^{n+1}$. Thus $M_{l}$ is a direct summand of $I^{j}$ with $0\leq j\leq n$. This shows that $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq{\rm{add}}(\oplus_{i=0}^{n}I^{i}).$
Theorem 1.10 (Miyachi [Mi00, Coro llary 1.3]).
Let
$0\longrightarrow Y_{n}\longrightarrow\cdots\longrightarrow Y_{1}\longrightarrow Y_{0}\longrightarrow X\longrightarrow 0$
be an exact sequence of $\Gamma$-modules. If
$0\longrightarrow Y_{i}\longrightarrow I^{-i,0}\longrightarrow I^{-i,1}\longrightarrow\cdots$
is an injective resolution of $Y_{i}$ ($0\leq i\leq n$), then $X$ has the injective resolution
$0\longrightarrow X\longrightarrow T\longrightarrow\bigoplus_{r+s=1}I^{r,s}\longrightarrow\bigoplus_{r+s=2}I^{r,s}\longrightarrow\cdots\longrightarrow\bigoplus_{r+s=n}I^{r,s}\longrightarrow\cdots$
where $T$ is a direct summand of $\bigoplus_{r+s=0}I^{r.s}$.
1.10. Infinitely Generated Modules
For later use, we record some facts about infinitely generated modules over an Artin algebra.
Facts 1.11.
Let $\Gamma$ is an Artin algebra.
  • (a)
    ([CE56, p. 17, Exercise 8]) Since $\Gamma$ is left Noetherian, arbitrary direct sums of injective (left) modules are injective modules, see also [Ch60, Proposition 4.1];
  • (b)
    ([KMT, Theorem 3.2]) Since $\Gamma$ is left Artinian, every injective module is a direct sum of indecomposable injective modules; see also Matlis [Ma58, Theorem 2.5];
  • (c)
    (Chase [Ch60, Theorem 3.3]) Since $\Gamma$ is left prefect and right coherent, direct products of projective modules are projective modules;
  • (d)
    (Müller [Mu70, Theorem 3]) Since $\Gamma$ is a semi-perfect ring, each projective module is a direct sum of finitely generated indecomposable projective modules.
@@ 1.11. Gorenstein Projective Modules
such that $M\simeq\mathrm{Im}(P_{0}\longrightarrow P^{0})$ and such that ${\rm{Hom}}_{\Gamma}(-,Q)$ leaves the sequence ${\bf P}$ exact whenever $Q$ is a projective module.
The class of all Gorenstein projective modules is denoted by ${\mathrm{GProj}}(\Gamma)$ and ${\mathrm{Gproj}}(\Gamma)$ denotes the subcategory of finitely generated Gorenstein projective modules; see [EJ11, Definition 10.2.1].
Gorenstein injective modules are defined dually. Denote by ${\mathrm{GInj}}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$ consisting of all Gorenstein injective $\Gamma$-modules and by ${\mathrm{Ginj}}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated Gorenstein injective $\Gamma$-modules; see [EJ95, Definition 2.1] or [EJ11, Definition 10.1.1].
The Gorenstein projective dimension of a $\Gamma$-module $M$ is at most $n\ (n\in\mathbb{N}_{0})$ if and only if $M$ has a Gorenstein projective resolution of length $n$, written as ${\rm{Gpd}}({}_{\Gamma}M)\leq n$. By [Ho04a, Theorem 2.10], ${\rm{Gpd}}({}_{\Gamma}M)\leq n$ if and only if there exists an exact sequence
$0\longrightarrow P_{n}\longrightarrow\cdots\longrightarrow P_{1}\longrightarrow G\longrightarrow M\longrightarrow 0,$
where $P_{n},\dots,P_{1}$ are projective modules and $G$ is a Gorenstein projective modules. Similarly, one can define the Gorenstein injective dimension ${\rm{Gid}}({}_{\Gamma}M)$ of $M$ [Ho04a, Definition 2.8]. Note that if ${\rm{pd}}({}_{\Gamma}M)$ is finite, then ${\rm{Gpd}}({}_{\Gamma}M)={\rm{pd}}({}_{\Gamma}M)$ [Ho04a, Proposition 2.27]. For every non-negative integer $m$, we denote by ${\mathrm{GProj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$, consisting of all $\Gamma$-modules of Gorenstein projective dimension at most $m$ and by ${\mathrm{Gproj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$, consisting of all finitely generated $\Gamma$-modules of Gorenstein projective dimension at most $m$.
Like the classical case, the supremum of Gorenstein projective dimensions of all left $\Gamma$-modules is called left Gorenstein global dimension of $\Gamma$ and is denoted by ${\rm{Ggldim}}(\Gamma)$. By [BM10, Theorem 1.1], the left Gorenstein global dimension of $\Gamma$ is also the supremum of Gorenstein injective dimensions of all left $\Gamma$-modules. By [EJ11, Theorem 12.3.1], there is also no difference between left and right Gorenstein global dimensions of $\Gamma$ (as every Artin algebra is Noetherian).
Also, the Gorenstein global dimension of $\Gamma$ is the supremum of Gorenstein projective dimensions of all finitely generated $\Gamma$-modules; see [Ch10, Theorem 3.2.8]. Moreover, ${\rm{Ggldim}}(\Gamma)\leq n$ if and only if $\Gamma$ is an $n$-Iwanaga-Gorenstein algebra (i.e. ${\rm{id}}({}_{\Gamma}\Gamma)\leq n$ and ${\rm{id}}(\Gamma_{\Gamma})\leq n$). In fact, in this case, we have ${\rm{Ggldim}}(\Gamma)={\rm{id}}({}_{\Gamma}\Gamma)={\rm{id}}(\Gamma_{\Gamma})$; see [EJ11, Corollary 12.3.2] or [Ch10, Corollary 3.2.6].
For any left $\Gamma$-module $M$ with finite projective dimension, ${\rm{Gpd}}({}_{\Gamma}M)\leq n$ if and only if ${\rm{Ext}}^{i}_{\Gamma}(M,\Gamma$ )=0)=0) = 0 for all $i>n$; see [Ho04a, Theorem 2.20]. Also, the Gorenstein injective dimension of $M$, ${\rm{Gid}}({}_{\Gamma}M)$, equals the usual injective dimension of $M$, i.e. if $M$ is a $\Gamma$-module with ${\rm{pd}}({}_{\Gamma}M)<\infty$, then ${\rm{Gid}}({}_{\Gamma}M)={\rm{id}}({}_{\Gamma}M)$; see [Ho04b, Theorem 2.1]. This result has a dual counterpart and if $M$ is a $\Gamma$-module with ${\rm{id}}({}_{\Gamma}M)<\infty$, then ${\rm{Gpd}}({}_{\Gamma}M)={\rm{pd}}({}_{\Gamma}M)$; see [Ho04b, Theorem 2.2].
@@ 2. Higher Auslander-Gorenstein Algebras
In this section, we will study higher Auslander-Gorenstein algebras introduced by Iyama and Solberg in [IS18] as a generalization of higher Auslander algebras and give a new characterization of them in terms of the abelianness property of the category of modules of Gorenstein projective dimension less than $n$. Of course, the results can be considered as a higher generalization of Auslander-Tachikawa theorem.
@@
Let $n$ be a positive integer. Recall that an Artin algebra $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra if ${\rm{id}}({}_{\Gamma}\Gamma)\leq n+1\leq\rm{domdim}(\Gamma)$ [IS18, Definition 1.1]. $\Gamma$ is also called a higher Auslander-Gorenstein algebra, if $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra for a certain $n$.
By [AR94, Corollary 5.5], these algebras are Iwanaga-Gorenstein algebras. It is also easily checked that any $n$-minimal Auslander-Gorenstein algebra $\Gamma$ is either selfinjective or satisfies ${\rm{id}}({}_{\Gamma}\Gamma)=n+1={\rm{domdim}}(\Gamma)$; see [IS18, Proposition 4.1].
Notice that, as mentioned in Subsection 1.11, for Gorenstein algebras, there is no difference between Gorenstein global dimention and injective dimension of $\Gamma$, i.e. $\mathrm{Ggldim}(\Gamma)={\rm{id}}({}_{\Gamma}\Gamma)$, and so $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra if and only if ${\rm{Ggldim}}(\Gamma)\leq n+1\leq\rm{domdim}(\Gamma)$.
@@
such that $E^{0},E^{1}\in{{\rm{add}}(\mathcal{Q})}$. Since $\mathcal{Q}$ is also projective, $E^{0}$ and $E^{1}$ are also projective. Let $X^{\prime}$ be the cokernel of the morphism $E^{0}\longrightarrow E^{1}$. Since ${\rm{Ggldim}}(\Gamma)\leq n+1$, ${\rm{Gpd}}({}_{\Gamma}X^{\prime})\leq n+1$, so by [Ho04a, Theorem 2.20] $,{\rm{Gpd}}({}_{\Gamma}X)\leq n-1$, i.e. $X\in{\mathrm{Gproj}}^{\leq n-1}(\Gamma)$. This shows that ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)\supseteq\mathrm{sub}^{2}(\mathcal{Q})$.
@@
we have ${\rm{Gpd}}({}_{\Gamma}K)\leq 1$ and by [Ho04a, Theorem 2.10], there exists an exact sequence
@@
such that $G$ is a finitely generated Gorenstein projective module and $P^{\prime}$ is a finitely generated projective $\Gamma$-module. Since ${\rm{domdim}}(\Gamma)\geq n+1\geq 2$, both $\mathbf{E}(G)$ and $\mathbf{E}(\mathbf{E}(P^{\prime})/P^{\prime})$ are projective, so by Theorem 1.10, it follows that $\mathbf{E}(K)$ is a projective $\Gamma$-module as well. Hence, $X$ fits into
@@
Now assume that $X$ is a $\Gamma$-module such that ${\rm{Gpd}}({}_{\Gamma}X)\leq n-1$. By [Ho04a, Theorem 2.10], there exists an exact sequence
@@
where $P_{i}$ is a finitely generated projective module for every $11\leq i\leq n-$ 1 and $G$ is a finitely generated Gorenstein projective module. Since $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra, ${\rm{Ggldim}}(\Gamma)\leq n+1\leq{\rm{domdim}}(\Gamma)$ and so every projective $\Gamma$-module $P_{i}$ has the minimal injective resolution
@@
such that $J^{-i,k}\in\mathrm{proj}(\Gamma)\ \cap\ \mathrm{inj}(\Gamma)$, for every $k=0,\dots,n$. Applying Theorem 1.10 to the exact sequence (2.2) together with (2.1) and (2.3) then immediately gives the required result.
@@
Let $\Gamma$ be an Artin algebra. If a full additive category $\mathcal{C}$ of ${\rm{{mod}}}(\Gamma)$ has kernels and contains $\Gamma$, then the kernels in $\mathcal{C}$ coincident with the kernels in ${\rm{{mod}}}(\Gamma)$ and the inclusion functor from $\mathcal{C}$ to ${\rm{{mod}}}(\Gamma)$ is left exact.
@@
[…]
@@
Let $\Gamma$ be an Artin algebra and $n\geq 1,0\leq k\leq n+1$. The following statements are equivalents:
@@
$(a)\Rightarrow(b)$: We prove by induction on $k$. For $k=0$, this is trivial. For $k=1$, Let $X\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$ and $Y$ be a submodule of $X$. Form the short exact sequence $0\to Y\to X\to Z\to 0$. As ${\rm{Ggldim}}(\Gamma)\leq n+1$, ${\rm{Gpd}}({}_{\Gamma}Z)\leq n+1$, and combining with ${\rm{Gpd}}({}_{\Gamma}X)\leq n$, we get ${\rm{Gpd}}({}_{\Gamma}Y)\leq n$ by [Ho04a, Theorem 2.20].
@@
where the last inclusion holds by inductive assumption. Combining with ${\rm{Gpd}}({}_{\Gamma}X^{0})\leq n-k$, we get ${\rm{Gpd}}({}_{\Gamma}Y)\leq n-k$ by[Ho04a, Theorem 2.20].
$(b)\Rightarrow(a)$: Let $X\in{\rm{{mod}}}(\Gamma)$ and $k\geq 1$. Given a presentation of $X$ by finitely generated projective modules:
@@
Before giving the proof, we would like to mention that torsion pairs play an important rule in the proof of Tachikawa’s theorem ([Ta74, Theorem 1]), Iyama’s theorem ([Iy04, Theorem 3.3]), and Kong’s theorem ([Ko14, Theorem 2.1]). We also follow this procedure and prove the result in four steps.
@@
Since ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)$ is abelian, it has kernels which, by Lemma 2.2, coincident with the kernels in ${\rm{{mod}}}(\Gamma)$. Hence, ${\rm{Ggldim}}(\Gamma)\leq n+1$ follows from Lemma 2.3.

### v2

@@ 2.5. Cotorsion Pairs
Let now $\mathcal{X}$ and $\mathcal{Y}$ be full subcategories of the abelian category $\mathcal{A}$. Recall that a pair $(\mathcal{X},\mathcal{Y})$ is called a cotorsion pair (or cotorsion theory) if $\mathcal{X}={}^{\perp_{1}}\mathcal{Y}$ and $\mathcal{Y}=\mathcal{X}^{\perp_{1}}$, see for instance, [18, Definition 7.1.2]. In this case, the class $\mathcal{X}$ is called a cotorsion class and the class $\mathcal{Y}$ is called a cotorsion-free class.
@@ 2.6. Projectivization and Injectivization
restricts to an equivalence $\mathrm{add}(M)\simeq\mathrm{proj}(\Gamma)$; see, for instance, [5, Proposition II.2.1 (c)].
@@
The first statement is well known; see, for instance, [5, Proposition II.2.5].
For the second statement, since $P$ is finitely generated, $\mathrm{Hom}_{\Lambda}(P,-)$ commutes with direct sums, thus it establishes an equivalence between $\mathrm{Add}(P)$ and ${\mathrm{Proj}}(\Gamma)$. Then the proof proceeds as that of [5, Proposition II.2.5]. ∎
@@
(a) This is the first lemma of [3, Chapter III, Section 4, Page 48].
@@
The two functors $\mathbf{D}\mathrm{Hom}_{\Gamma}(-,\mathcal{Q})$ and $\mathrm{Hom}_{\Gamma}(I,-)$ are exact, in particular, right exact, so by Watts’ theorem (see, for example, [40, Corollary 5.47]), they are isomorphic to a tensor functor. It suffices to check that $\mathbf{D}\mathrm{Hom}_{\Gamma}(\Gamma,\mathcal{Q})\cong\mathbf{D}\mathcal{Q}\cong\mathrm{Hom}_{\Gamma}(I,\Gamma)$. In fact, since $\mathbf{D}\mathcal{Q}$ is finitely generated projective, its double $\Gamma$-dual $\mathrm{Hom}_{\Gamma}(\mathrm{Hom}_{\Gamma^{{\rm{op}}}}(\mathbf{D}\mathcal{Q},\Gamma),\Gamma)=\mathrm{Hom}_{\Gamma}(I,\Gamma)$ is isomorphic to itself.
@@
commutes with direct sums, so the equivalence $\mathrm{add}(\mathcal{Q})\simeq\mathrm{inj}(\Lambda)$ extends to $\mathrm{Add}(\mathcal{Q})\simeq{\mathrm{Inj}}(\Lambda)$. As in the proof of the first lemma of [3, Chapter III, Section 4], the latter equivalence extends to another equivalence
@@
2.7. Iwanaga-Gorenstein algebras
Recall that an Artin algebra $\Gamma$ is called Iwanaga-Gorenstein if its injective dimensions both as a left and as a right $\Gamma$-module are finite, which are equal by Zaks [45, Lemma A] ( see also [18, Proposition 9.1.8]).
For Iwanaga-Gorenstein algebras of injective dimension $n+1$, we have the following proposition.
@@ 2.7. Iwanaga-Gorenstein algebras
$0\longrightarrow{{}_{\Gamma}\Gamma}\longrightarrow I^{0}\longrightarrow\cdots\longrightarrow I^{n}\longrightarrow I^{n+1}\to 0.$
@@
If $M\in\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)$, then $M=\bigoplus_{l=1}^{r}M_{l}$, where $M_{l}$ is an injective indecomposable $\Gamma$-module and of projective dimension at most $n$. By [23, Theorem 2], each $M_{l}$ is a direct summand of $I^{j}$ with $0\leq j\leq n+1$. But ${\rm{pd}}(M_{l})\leq n$, so by [24, Theorem 2], $M_{l}$ is not a direct summand of $I^{n+1}$. Thus $M_{l}$ is a direct summand of $I^{j}$ with $0\leq j\leq n$. This shows that $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq{\rm{add}}(\oplus_{i=0}^{n}I^{i}).$
@@ 2.8. Gorenstein Projective Modules
such that $M\simeq\mathrm{Im}(P_{0}\longrightarrow P^{0})$ and such that ${\rm{Hom}}_{\Gamma}(-,Q)$ leaves the sequence $\bf P$} exact whenever $Q$ is a projective module; see [18, Definition 10.2.1].
The class of all Gorenstein projective modules is denoted by ${\mathrm{GProj}}(\Gamma)$ and ${\mathrm{Gproj}}(\Gamma)$ denotes the subcategory of finitely generated Gorenstein projective modules.
The Gorenstein projective dimension of a $\Gamma$-module $M$ is at most $n\ (n\in\mathbb{N}_{0})$ if $M$ has a Gorenstein projective resolution of length $n$, written as ${\rm{Gpd}}({}_{\Gamma}M)\leq n$. For every non-negative integer $m$, we denote by ${\mathrm{GProj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$ consisting of all $\Gamma$-modules of Gorenstein projective dimension at most $m$ and by ${\mathrm{Gproj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated $\Gamma$-modules of Gorenstein projective dimension at most $m$.
Like the classical case, the supremum of Gorenstein projective dimensions of all left $\Gamma$-modules is called left Gorenstein global dimension of $\Gamma$ and is denoted by ${\rm{Ggldim}}(\Gamma)$. Also, ${\rm{Ggldim}}(\Gamma)\leq n$ if and only if $\Gamma$ is an $n$-Iwanaga-Gorenstein algebra (i.e. ${\rm{id}}({}_{\Gamma}\Gamma)\leq n$ and ${\rm{id}}(\Gamma_{\Gamma})\leq n$). In fact, in this case, we have ${\rm{Ggldim}}(\Gamma)={\rm{id}}({}_{\Gamma}\Gamma)={\rm{id}}(\Gamma_{\Gamma})$; see [18, Corollary 12.3.2] or [14, Corollary 3.2.6].
@@ 3. Higher Auslander-Gorenstein Algebras
In this section, we will study higher Auslander-Gorenstein algebras introduced by Iyama and Solberg in [31] as a generalization of higher Auslander algebras and give a new characterization of them in terms of the abelianness property of the category of modules of Gorenstein projective dimension less than $n$. Of course, the results can be considered as a higher generalization of Auslander-Tachikawa theorem.
@@
Let $n$ be a positive integer. Recall that an Artin algebra $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra if ${\rm{id}}({}_{\Gamma}\Gamma)\leq n+1\leq\rm{domdim}(\Gamma)$ [31, Definition 1.1]. $\Gamma$ is also called a higher Auslander-Gorenstein algebra, if $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra for a certain $n$.
By [4, Corollary 5.5], these algebras are Iwanaga-Gorenstein algebras. It is also easily checked that any $n$-minimal Auslander-Gorenstein algebra $\Gamma$ is either selfinjective or satisfies ${\rm{id}}({}_{\Gamma}\Gamma)=n+1={\rm{domdim}}(\Gamma)$; see [31, Proposition 4.1].
Notice that, as mentioned in Subsection 2.8, for Iwanaga-Gorenstein algebras, there is no difference between Gorenstein global dimension and injective dimension of $\Gamma$, i.e. $\mathrm{Ggldim}(\Gamma)={\rm{id}}({}_{\Gamma}\Gamma)$, and so $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra if and only if ${\rm{Ggldim}}(\Gamma)\leq n+1\leq\rm{domdim}(\Gamma)$.
@@
such that $E^{0},E^{1}\in{{\rm{add}}(\mathcal{Q})}$. Since $\mathcal{Q}$ is also projective, $E^{0}$ and $E^{1}$ are also projective. Let $X^{\prime}$ be the cokernel of the morphism $E^{0}\longrightarrow E^{1}$. Since ${\rm{Ggldim}}(\Gamma)\leq n+1$, ${\rm{Gpd}}({}_{\Gamma}X^{\prime})\leq n+1$, so by [20, Theorem 2.20] ${\rm{Gpd}}({}_{\Gamma}X)\leq n-1,$ i.e. $X\in{\mathrm{Gproj}}^{\leq n-1}(\Gamma)$. This shows that ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)\supseteq\mathrm{sub}^{2}(\mathcal{Q})$.
@@
we have ${\rm{Gpd}}({}_{\Gamma}K)\leq 1$ and by [20, Theorem 2.10], there exists an exact sequence
@@
such that $G$ is a finitely generated Gorenstein projective module and $P^{\prime}$ is a finitely generated projective $\Gamma$-module. Since ${\rm{domdim}}(\Gamma)\geq n+1\geq 2$, both $\mathbf{E}(G)$ and $\mathbf{E}(\mathbf{E}(P^{\prime})/P^{\prime})$ are projective, so by [38, Corollary 1.3], it follows that $\mathbf{E}(K)$ is a projective $\Gamma$-module as well. Hence, $X$ fits into
@@
Now assume that $X$ is a $\Gamma$-module such that ${\rm{Gpd}}({}_{\Gamma}X)\leq n-1$. By [20, Theorem 2.10], there exists an exact sequence
@@
where $P_{i}$ is a finitely generated projective module for every $1\leq i\leq n-1$ and $G$ is a finitely generated Gorenstein projective module. Since $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra, ${\rm{Ggldim}}(\Gamma)\leq n+1\leq{\rm{domdim}}(\Gamma)$ and so every projective $\Gamma$-module $P_{i}$ has the minimal injective resolution
@@
such that $J^{-i,k}\in\mathrm{proj}(\Gamma)\ \cap\ \mathrm{inj}(\Gamma)$, for every $k=0,\dots,n$. Applying [38, Corollary 1.3] to the exact sequence (3.2) together with (3.1) and (3.3) then immediately gives the required result.
@@
Let $\Gamma$ be an Artin algebra. If a full additive category $\mathcal{C}$ of ${\rm{{mod}}}(\Gamma)$ has kernels and contains $\Gamma$, then the kernels in $\mathcal{C}$ coincides with the kernels in ${\rm{{mod}}}(\Gamma)$ and the inclusion functor from $\mathcal{C}$ to ${\rm{{mod}}}(\Gamma)$ is left exact.
@@
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\rm{Hom}}_{\Gamma}(\Gamma,K^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${{\rm{Hom}}_{\Gamma}(\Gamma,X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${{\rm{Hom}}_{\Gamma}(\Gamma,Y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{k^{\prime}}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Y}$
@@
Let $\Gamma$ be an Artin algebra and $n\geq 1,0\leq k\leq n+1$. The following statements are equivalent:
@@
$(a)\Rightarrow(b)$. We prove by induction on $k$. For $k=0$, this is trivial. For $k=1$, let $X\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$ and $Y$ be a submodule of $X$. Form the short exact sequence $0\to Y\to X\to Z\to 0$. As ${\rm{Ggldim}}(\Gamma)\leq n+1$, ${\rm{Gpd}}({}_{\Gamma}Z)\leq n+1$, and combining with ${\rm{Gpd}}({}_{\Gamma}X)\leq n$, we get ${\rm{Gpd}}({}_{\Gamma}Y)\leq n$ by [20, Theorem 2.20].
@@
where the last inclusion holds by inductive assumption. Combining with ${\rm{Gpd}}({}_{\Gamma}X^{0})\leq n-k$, we get ${\rm{Gpd}}({}_{\Gamma}Y)\leq n-k$ by [20, Theorem 2.20].
$(b)\Rightarrow(a)$. Let $X\in{\rm{{mod}}}(\Gamma)$ and $k\geq 1$. Given a resolution of $X$ by finitely generated projective modules:
@@
Before giving the proof, we would like to mention that torsion pairs play an important rule in the proof of Tachikawa theorem ([44, Theorem 1]), Iyama’s theorem ([25, Theorem 3.3]), and Kong’s theorem ([33, Theorem 2.1]). We also follow this procedure and prove the result in three steps.
@@
Since ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)$ is abelian, it has kernels which, by Lemma 3.2, coincides with the kernels in ${\rm{{mod}}}(\Gamma)$. Hence, ${\rm{Ggldim}}(\Gamma)\leq n+1$ follows from Lemma 3.3.